by using Matlab R20 16a

I need the code and the graph

* Write a program to plot the derivative of function ?=?.?^−?^?
Answer #1

MATLAB R2018b-academic use HOME PLOTS ② arch Documentation Sign In O Preferences Set Path Parallel ▼ New Variable Analyze Code Find Files EZ Open Variable ▼ Run and Time Request Support Simulink Layout New New New Open Compare ImportSave Script Live Script Favorites Add-Ons Help Data Workspace Ciear Workspace Clear Commands ▼ Learn WATLAB Current Folder Command Window Name New to MATLAB? See resources for Gatting Started 田□ 10-9 困) 13-9 E numerical lab syms ti x-t*exp(-t2) xdot = diff (x, t, 1) t -0:0.01:3; x-= subs(x, t, t-) ; xdot-subs (xdot, t,t pp tut 困 -i project 1.jp9 1a.jpg b.jpg cjpg 1d.jpg 1d2.jpg 1d3jpg x = t*exp (-t**2); Error: Invalid use of operator syms ti xtexp(-t*t) xdot = diff (x, t, 1) t -0:0.01:3; x-= subs (x, t, t-) ; xdot-subs (xdot, t,t 1di2 Details Workspace Name ▲ Value 1xl sym x301 double 1x1 sym 1x301 sym 1xl sym 1x301 sym xdot xdot xdot exp (-t^2) -2*t2exp (-t^2) O Type here to search 1/202019 6MATLAB R2018b-academic use Figure 1 HOME PLOTS ② arch Documentation Sign In File Edit View Insert Tools Desktop Window Help Find Files 笛1 必風ロ困 E Request Support New New New Open Compare ImportSave Script Live Script Help Data Workspace Learn WATLAB Current Folder Name- New to 田□ 10-9 困) 13-9 3 x- xdot 0.5 numerical lab x E xdot 困 -i project 1.jp9 1a.jpg b.jpg cjpg 1d.jpg 1d2.jpg 1d3jpg plot 1di2 0.5 Details Workspace Name ▲ xdot 0.5 2.5 Value 1xl sym x301 double 1x1 sym 1x301 sym 1xl sym 1x301 sym xdot-subs (xdot, t,t xdot xdot xdot exp (-t^2) -2*t2exp (-t^2) O Type here to search 1/202019 6

please rate it up thanks :)